# The KSBA moduli space of stable log Calabi–Yau surfaces

### 23 Feb 2024 v1

@@ 1.4. Related work
Mirrors to log smooth maximal degenerations to (log) Calabi–Yau varieties are constructed by Gross–Siebert in [47]. As the maximal degenerations of polarized log Calabi–Yau surfaces we consider in this paper are typically not log smooth, the construction of [47] does not apply directly to produce the semistable mirror $\mathcal{X}\rightarrow\Delta$. Also, the mirror construction of Gross–Hacking–Keel [40] concerns a single “smooth” log Calabi–Yau surface, without maximal degenerations, and hence it does not apply either. To address this issue, we give a direct construction by deformation theory of the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. If we restrict attention to smooth log Calabi–Yau surfaces, we expect the affinization $\mathcal{\overline{X}}_{\mathscr{P}}$ of our semistable mirror $\mathcal{X}_{\mathscr{P}}$ to be deformation equivalent to the mirror of [40]. However, the existence of a crepant resolution in that setup is not immediate. Moreover, in order to eventually relate the geometry of the “double mirror” to the original polarized log Calabi–Yau surface, our original semistable mirror construction in our paper plays a crucial role.
@@
To construct the intrinsic mirror to $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we use a version of the Gross–Siebert mirror construction [47, 48], which is possible since $\mathcal{X}\rightarrow\Delta$ is log smooth. In [49], the non-Archimedean mirror construction developed by Keel–Yu [63] is used instead to verify the Hacking–Keel–Yu conjecture for $6$ stable log Calabi–Yau surfaces $(Y,D+\epsilon\,C)$, namely when $Y$ is a del Pezzo surfaces of degree $n$, for $1\leq n\leq 6$, endowed with the line bundle $L=-K_{Y}$, and $D$ a cycle of $n$ $(-1)$-curves. These surfaces are in a sense “self-mirror”, for example, the dimension of the complex moduli equals the dimension of the Kähler moduli, and due to this coincidence one can conveniently take the mirror $\mathcal{X}$ to any of these del Pezzo’s $Y$ as the canonical bundle $K_{Y}$. In this paper, to prove the Hacking–Keel–Yu conjecture for all log Calabi–Yau surfaces, we investigate double mirrors by studying intrinsic mirrors to our semistable mirrors, applying (a version of) the construction of [47] rather than the non-Archimedean mirror construction of [63]. To be able to use the non-Archimedean mirror construction in this generality, it would first be necessary to have an extension of the results of [63] over a field with nontrivial valuation, which is not yet available. [^1] After the second author presented results of this paper at the Homological Mirror Symmetry conference on January 28th 2024 at IMSA/Miami, we were informed that Logan White is currently working on such an extension in his PhD under the guidance of Sean Keel.
@@
Finally, the ongoing work of Gross–Hacking–Keel–Siebert [43] on describing compactifications of moduli spaces of K3 surfaces using intrinsic mirror symmetry is also closely related, as open Kulikov degenerations introduced in our paper are non-compact analogs of the Type III Kulikov degenerations of K3 surfaces – see also [53] and [54] in this context.
@@ 2.1. Log Calabi–Yau surfaces
In this paper, a log Calabi–Yau surface is a log canonical (lc) pair $(Y,D)$, as in [67, Definition 2.34], with $Y$ projective of dimension $2$, $D$ a reduced effective divisor, and $K_{Y}+D$ trivial (and in particular Cartier). Throughout this paper we assume that all log Calabi–Yau surfaces are maximal, in the sense that $D$ is non-empty and singular. The following proposition classifies singularities of log Calabi–Yau surfaces.
@@
We will consider polarized log Calabi–Yau surfaces $(Y,D,L)$, given by a log Calabi–Yau surface $(Y,D)$ endowed with an ample line bundle $L$. In order to define below the notion of a toric model for a polarized log Calabi–Yau surface, we first define interior and corner blow-ups of log Calabi–Yau surfaces.
@@
where $g$ is a corner blow-up, $h$ is a composition of interior blow-ups, $(\overline{Y},\overline{D})$ is a toric log Calabi–Yau surface, and such that the line bundle $\overline{L}:=h_{\star}g^{\star}L$ is ample on $\overline{Y}$.
Every Calabi-Yau surface admits a toric model by [40, Lemma 1.3] applied to the minimal resolution given by Corollary 2.2. We will use particularly good toric models, defined as follows.
@@
    $h$ is a composition of interior blow-ups at distinct points of $\overline{D}$,
@@ 2.2. Symington polytopes
Associated to a polarized log Calabi–Yau surface $(Y,D,L)$, there is a so called Symington “polytope” $P$, given by an integral affine manifold with singularities, which corresponds to the base of an almost toric fibration $Y\to P$ [90, §$2.2$]. Engel and Friedman showed in [31] that one can make nice choices of Symington polytopes, so that they are obtained from a toric momentum polytope by cutting out non-overlapping triangles with one edge adjacent to the boundary. In this section we review this construction, following [31, §$5.1$].
@@
We describe the construction of a Symington polytope for a polarized log Calabi–Yau surface $(Y,D,L)$ endowed with a good toric model $(\overline{Y},\overline{D},\overline{L})$ as in Definition 2.7. We denote by $\overline{P}$ the toric momentum polytope associated to the ample line bundle $\overline{L}$ on the toric surface $(\overline{Y},\overline{D})$. The edges $(\overline{P}_{i})_{1\leq i\leq n}$ of the boundary $\partial\overline{P}$ of $\overline{P}$ are naturally in one to one correspondence with the toric irreducible components $(\overline{D}_{i})_{1\leq i\leq n}$ of $\overline{D}$. Moreover, for every $1\leq i\leq n$, the integral length $\ell(\overline{P}_{i})$ of the edge $\overline{P}_{i}$ is given by the intersection number $\overline{L}\cdot\overline{D}_{i}$.
As in Definition 2.7, we denote the exceptional curves of the blow-up $h:\widetilde{Y}\to\overline{Y}$ contracted to $\overline{D}_{i}$ by $E_{ij}$, for $1\leq j\leq N_{i}$. By Definition 2.7 of a good toric model, there exists a decomposition
@@
We refer to the vertex $p_{ij}$ of $\Delta_{ij}$ as the interior vertex of $\Delta_{ij}$, and to the edges of $\Delta_{ij}$ adjacent to $p_{ij}$ as the interior edges of $\Delta_{ij}$. Note that, sometimes the interior vertex of a triangle $\Delta_{ij}$ may coincide with the interior vertex of another surgery triangle, leading to more complex singularities on $Y$ than focus-focus singularities, when necessary in the following sections we will use appropriate perturbations, separating each singularity into a focus-focus one.
The following result shows that the monodromy invariant lines have always a non-trivial intersection with the Symington polytope $P$, that is, this intersection cannot be reduced to the single point $p_{ij}$.
@@
For every $1\leq i\leq n$, $1\leq j\leq N_{i}$, the intersection of the monodromy invariant line $L_{ij}$ with $P$ is an interval of positive length.
@@ 2.3. Good polyhedral decompositions
$\left\{(0,0),\left(\frac{a}{d},\frac{b}{d}\right),\left(\frac{a}{d},\frac{c}{d}\right)\right\}$ (2.4)
@@ 2.4. Maximal degenerations
One starts with the Mumford toric degeneration
$(\overline{\mathcal{Y}}_{\varphi},\overline{\mathcal{D}}_{\varphi},\overline{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$
of the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ induced by the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$ [38, Example 3.6]. The intersection complex of the central fiber $(\overline{Y}_{\varphi,0},\overline{D}_{\varphi,0},\overline{L}_{\varphi,0})$ is the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$: the irreducible components of the central fiber are the polarized toric surfaces $Y^{f}$ with momentum polytopes the 2-dimensional faces $f$ of $\overline{\mathscr{P}}$ and they are glued along toric divisors $Y^{e}$ corresponding to the edges $e$ of $\overline{\mathscr{P}}$.
For every $1\leq i\leq n$ and $1\leq j\leq N_{i}$, we pick a section $s_{ij}$ of
$(\overline{\mathcal{Y}}_{\varphi},\overline{\mathcal{D}}_{\varphi},\overline{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}\,,$
whose image in the general fiber is contained in the interior of the irreducible component $\overline{D}_{i}$ of $\overline{D}$, and whose image in the central fiber is a point $x_{ij}$ in the interior of the divisor $Y^{\overline{P}_{ij}}$ corresponding to the edge $\overline{P}_{ij}$ of $\overline{\mathscr{P}}$ in the boundary of $\overline{P}$, as in Construction 2.10.
@@
where $\mathcal{E}_{ij}$ are the exceptional divisors. Comparing with Equation (2.2), it follows that the general fiber of
$(\widetilde{\mathcal{Y}}_{\varphi},\widetilde{\mathcal{D}}_{\varphi},\widetilde{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}\,$
is deformation equivalent to $(\widetilde{Y},\widetilde{D},\widetilde{L})$. While $\widetilde{L}$ is a nef line bundle, the restriction ~$\widetilde{\mathcal{L}}_{\varphi,0}$ of ~$\widetilde{\mathcal{L}}_{\varphi}$ to the central fiber $(\widetilde{\mathcal{Y}}_{\varphi,0},\widetilde{\mathcal{D}}_{\varphi,0})$ is not nef if there exists $i$ and $j$ with $m_{ij}>1$.
Indeed, if $m_{ij}>1$, then the exceptional divisor $\mathcal{E}_{ij,0}$ on the central fiber is obtained by blowing-up a point on the divisor $Y^{\overline{P}_{ij}}$, which is contained in the toric surface $Y^{\Delta_{ij,m_{ij}}}_{\overline{\mathscr{P}}}$ with momentum polytope given by the trapezoid $\Delta_{ij,m_{ij}}$ as in Equation (2.5). As $\Delta_{ij,m_{ij}}$ is a trapezoid of size $1$, it follows that $X^{\Delta_{ij,m_{ij}}}_{\overline{\mathscr{P}}}$ is a $\mathbb{P}^{1}$-fibration whose fiber class $F$ satisfies ¯$\overline{\mathcal{L}}_{\varphi,0}\cdot F=1$. After the blow-up, the strict transform of the fiber class passing through the blown-up point is a $(-1)$-curve $C_{ij}$ of class $F-\mathcal{E}_{ij,0}$, and so we have
$\widetilde{L}_{\varphi,0}\cdot C_{ij}=1-m_{ij}\,,$
which is negative if $m_{ij}>1$.
@@
In order to obtain a nef line bundle, we first replace the 3-fold $\widetilde{Y}_{\varphi}$ by its flop along the $(-1,-1)$-curve $C_{ij}$. On the special fiber, this means blowing up the component corresponding to $\Delta_{ij,m_{ij}-1}$ instead of $\Delta_{ij,m_{ij}-1}$. If $m_{ij}>2$, then the line bundle is not nef either, because the intersection with the strict transform of the $\mathbb{P}^{1}$-fiber containing the blown-up point is $2-m_{ij}$, and one can flop this curve – see Figure 2.3. In general, after $m_{ij}-1$ flops, we obtain a family
$(\widetilde{\mathcal{Y}}^{\prime}_{\varphi},\widetilde{\mathcal{D}}^{\prime}_{\varphi},\widetilde{\mathcal{L}}^{\prime}_{\varphi})\rightarrow\mathbb{A}^{1}\,,$
whose general fiber is still deformation equivalent to $(\widetilde{Y},\widetilde{D},\widetilde{L})$, but with now the line bundle ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ being relatively big and nef. By [34, Theorem 4.18], a big enough power of ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ is base point free, and so defines a relatively birational contraction
@@
$(\mathcal{Y}_{\varphi},\mathcal{D}_{\varphi},\mathcal{L}_{\varphi})\rightarrow\mathbb{A}^{1}$ (2.8)
@@
In the following sections, we describe the “mirror” to the degeneration in Equation (2.8), which we refer to the semistable mirror – as explained above, this degeneration is obtained from the data of the polarized log Calabi–Yau $(Y,D,L)$ together with the choice of the Symington polytope $P$ and a good polyhedral decomposition $\mathscr{P}$, as well as a choice of a compatible PL function $\varphi$, we will refer to the mirror of the degeneration in Equation (2.8), as the mirror to $(Y,D,L)$ together with $P$ and $\mathscr{P}$, suppressing the date of $\varphi$. While the algebro-geometric mirror construction of Gross-Siebert applies to particular types of degenerations, namely to so called “toric” degenerations in the framework of [46] and to so called “log smooth” degenerations in the framework of [47, 48], the degeneration in Equation (2.8) is neither toric nor log smooth. Hence, the mirror construction we provide in this paper is different than the so far known constructions.

### 30 Mar 2024 v2

@@ 1.4. Related work
Mirrors to log smooth maximal degenerations to (log) Calabi–Yau varieties are constructed by Gross–Siebert in [48]. As the maximal degenerations of polarized log Calabi–Yau surfaces we consider in this paper are typically not log smooth, the construction of [48] does not apply directly to produce the semistable mirror $\mathcal{X}\rightarrow\Delta$. Also, the mirror construction of Gross–Hacking–Keel [41] concerns a single “smooth” log Calabi–Yau surface, without maximal degenerations, and hence it does not apply either. To address this issue, we give a direct construction by deformation theory of the semistable mirror $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$. If we restrict attention to smooth log Calabi–Yau surfaces, we expect the affinization $\mathcal{\overline{X}}_{\mathscr{P}}$ of our semistable mirror $\mathcal{X}_{\mathscr{P}}$ to be deformation equivalent to the mirror of [41]. However, the existence of a crepant resolution in that setup is not immediate. Moreover, in order to eventually relate the geometry of the “double mirror” to the original polarized log Calabi–Yau surface, the semistable mirror construction introduced in our paper plays a crucial role.
@@
To construct the double mirror to $(Y,D,L)$, that is, the intrinsic mirror to $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$, we use a version of the Gross–Siebert mirror construction [48, 49], which is possible since $\mathcal{X}_{\mathscr{P}}\rightarrow\Delta$ is log smooth. In [50], the non-Archimedean mirror construction developed by Keel–Yu [64] is used instead to verify the Hacking–Keel–Yu conjecture for $6$ stable log Calabi–Yau surfaces $(Y,D+\epsilon\,C)$, namely when $Y$ is a del Pezzo surfaces of degree $n$, for $1\leq n\leq 6$, endowed with the line bundle $L=-K_{Y}$, and $D$ a cycle of $n$ $(-1)$-curves. These surfaces are in a sense “self-mirror”, for example, the dimension of the complex moduli equals the dimension of the Kähler moduli, and due to this coincidence one can conveniently take the mirror to any of these del Pezzo’s $Y$ as the canonical bundle $K_{Y}$. In this paper, to prove the Hacking–Keel–Yu conjecture for all log Calabi–Yau surfaces, we investigate double mirrors by studying intrinsic mirrors to our semistable mirrors, applying (a version of) the construction of [48] rather than the non-Archimedean mirror construction of [64]. To be able to use the non-Archimedean mirror construction in this generality, it would first be necessary to have an extension of the results of [64] over a field with nontrivial valuation, which is not yet available [^1] After the second author presented results of this paper at the Homological Mirror Symmetry conference on January 28th 2024 at IMSA/Miami, we were informed that Logan White is currently working on such an extension in his PhD under the guidance of Sean Keel..
@@
Finally, the ongoing work of Gross–Hacking–Keel–Siebert [44] on describing compactifications of moduli spaces of K3 surfaces using intrinsic mirror symmetry is also closely related, as open Kulikov degenerations introduced in our paper are non-compact analogues of the Type III Kulikov degenerations of K3 surfaces – see also [54] and [55] in this context.
@@ 2.1. Log Calabi–Yau surfaces
In this paper, a log Calabi–Yau surface is a log canonical (lc) pair $(Y,D)$ over $\mathbb{C}$, as in [67, Definition 2.34], with $Y$ projective of dimension $2$, $D$ a reduced effective divisor, and $K_{Y}+D$ trivial, in particular Cartier. Throughout this paper we assume that all log Calabi–Yau surfaces are maximal, in the sense that $D$ is non-empty and singular. The following proposition classifies singularities of log Calabi–Yau surfaces.
@@
We will consider polarized log Calabi–Yau surfaces $(Y,D,L)$, given by a log Calabi–Yau surface $(Y,D)$ endowed with an ample line bundle $L$. In what follows we describe toric models for such surfaces. For this, we first define first define interior and corner blow-ups.
@@
where $g$ is a corner blow-up, $h$ is a composition of interior blow-ups, $(\overline{Y},\overline{D})$ is a toric log Calabi–Yau surface, and $\overline{L}:=h_{\star}g^{\star}L$ is an ample line bundle on $\overline{Y}$.
Every Calabi-Yau surface admits a toric model by [41, Lemma 1.3] applied to the minimal resolution given by Corollary 2.2. We will use particularly good toric models, defined as follows.
@@
    $h$ is a composition of interior blow-ups at distinct points of $\overline{D}$.
@@ 2.2. Symington polytopes
Associated to a polarized log Calabi–Yau surface $(Y,D,L)$, there is a so called Symington “polytope” $P$, given by an integral affine manifold with singularities, which corresponds to the base of an almost toric fibration $Y\to P$ [91, §$2.2$]. Engel and Friedman showed in [31] that one can make nice choices of Symington polytopes, so that they are obtained from a toric momentum polytope by cutting out non-overlapping triangles with one edge adjacent to the boundary. In this section we review this construction, following [31, §$5.1$].
@@
We describe the construction of a Symington polytope for a polarized log Calabi–Yau surface $(Y,D,L)$ endowed with a good toric model $(\overline{Y},\overline{D},\overline{L})$ as in Definition 2.7. We denote by $\overline{P}$ the toric momentum polytope associated to the ample line bundle $\overline{L}$ on the toric surface $(\overline{Y},\overline{D})$. The edges $(\overline{P}_{i})_{1\leq i\leq n}$ of the boundary $\partial\overline{P}$ of $\overline{P}$ are naturally in one-to-one correspondence with the toric irreducible components $(\overline{D}_{i})_{1\leq i\leq n}$ of $\overline{D}$. Moreover, for every $1\leq i\leq n$, the integral length $\ell(\overline{P}_{i})$ of the edge $\overline{P}_{i}$ is given by the intersection number $\overline{L}\cdot\overline{D}_{i}$.
As in Definition 2.7, we denote the exceptional curves of the blow-up $h:\widetilde{Y}\to\overline{Y}$ contracted to $\overline{D}_{i}$ by $E_{ij}$, for $1\leq j\leq N_{i}$. By Definition 2.7, there exists a decomposition
@@
We refer to the vertex $p_{ij}$ of $\Delta_{ij}$ as the interior vertex of $\Delta_{ij}$, and to the edges of $\Delta_{ij}$ adjacent to $p_{ij}$ as the interior edges of $\Delta_{ij}$. Note that, sometimes the interior vertices of several triangles $\Delta_{ij}$ may coincide, leading to more complex singularities on $Y$ than focus-focus singularities.
The following result shows that the monodromy invariant lines have always a non-trivial intersection with the Symington polytope $P$.
@@
For every $1\leq i\leq n$, $1\leq j\leq N_{i}$, the intersection of the monodromy invariant line $L_{ij}$ with the interior of $P$ is an interval of positive length.
@@ 2.3. Good polyhedral decompositions
$\left\{(0,0),\left(\frac{a}{d},\frac{b}{d}\right),\left(\frac{a}{d},\frac{c}{d}\right)\right\}\,,$ (2.4)
@@ 2.4. Maximal degenerations
We start with the Mumford toric degeneration $(\overline{\mathcal{Y}}_{\varphi},\overline{\mathcal{D}}_{\varphi},\overline{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$ of the polarized toric surface $(\overline{Y},\overline{D},\overline{L})$ induced by the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$ [39, Example 3.6]. The intersection complex of the central fiber $(\overline{Y}_{\varphi,0},\overline{D}_{\varphi,0},\overline{L}_{\varphi,0})$ is the polyhedral decomposition $\overline{\mathscr{P}}$ of $\overline{P}$: the irreducible components of the central fiber are the polarized toric surfaces $Y^{f}$ with momentum polytopes the 2-dimensional faces $f$ of $\overline{\mathscr{P}}$ and they are glued along toric divisors $Y^{e}$ corresponding to the edges $e$ of $\overline{\mathscr{P}}$.
For every $1\leq i\leq n$ and $1\leq j\leq N_{i}$, we pick a section $s_{ij}$ of $(\overline{\mathcal{Y}}_{\varphi},\overline{\mathcal{D}}_{\varphi},\overline{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$, whose image in the general fiber is contained in the interior of the irreducible component $\overline{D}_{i}$ of $\overline{D}$, and whose image in the central fiber is a point $x_{ij}$ in the interior of the divisor $Y^{\overline{P}_{ij}}$ corresponding to the edge $\overline{P}_{ij}$ of $\overline{\mathscr{P}}$ in the boundary of $\overline{P}$, as in Construction 2.10.
@@
where $\mathcal{E}_{ij}$ are the exceptional divisors. Comparing with Equation (2.2), it follows that the general fiber of $(\widetilde{\mathcal{Y}}_{\varphi},\widetilde{\mathcal{D}}_{\varphi},\widetilde{\mathcal{L}}_{\varphi})\rightarrow\mathbb{A}^{1}$ is deformation equivalent to $(\widetilde{Y},\widetilde{D},\widetilde{L})$. While $\widetilde{L}$ is a nef line bundle, the restriction ~$\widetilde{\mathcal{L}}_{\varphi,0}$ of ~$\widetilde{\mathcal{L}}_{\varphi}$ to the central fiber $(\widetilde{\mathcal{Y}}_{\varphi,0},\widetilde{\mathcal{D}}_{\varphi,0})$ is not nef if there exists $i$ and $j$ with $m_{ij}>1$. Indeed, if $m_{ij}>1$, then the exceptional divisor $\mathcal{E}_{ij,0}$ on the central fiber is obtained by blowing-up a point on the divisor $Y^{\overline{P}_{ij}}$, which is contained in the toric surface $Y^{\Delta_{ij,m_{ij}}}_{\overline{\mathscr{P}}}$ with momentum polytope given by the trapezoid $\Delta_{ij,m_{ij}}$ as in Equation (2.5). As $\Delta_{ij,m_{ij}}$ is a trapezoid of size $1$, it follows that $X^{\Delta_{ij,m_{ij}}}_{\overline{\mathscr{P}}}$ is a $\mathbb{P}^{1}$-fibration whose fiber class $F$ satisfies ¯$\overline{\mathcal{L}}_{\varphi,0}\cdot F=1$. After the blow-up, the strict transform of the fiber class passing through the blown-up point is a $(-1)$-curve $C_{ij}$ of class $F-\mathcal{E}_{ij,0}$, and so we have $\widetilde{L}_{\varphi,0}\cdot C_{ij}=1-m_{ij}$, which is negative if $m_{ij}>1$.
@@
In order to obtain a nef line bundle, we first replace the 3-fold $\widetilde{Y}_{\varphi}$ by its flop along the $(-1,-1)$-curve $C_{ij}$. On the special fiber, this means blowing up the component corresponding to $\Delta_{ij,m_{ij}-1}$ instead of $\Delta_{ij,m_{ij}-1}$. If $m_{ij}>2$, then the line bundle is not nef either, because the intersection with the strict transform of the $\mathbb{P}^{1}$-fiber containing the blown-up point is $2-m_{ij}$, and one can flop this curve – see Figure 2.3. In general, after $m_{ij}-1$ flops, we obtain a family $(\widetilde{\mathcal{Y}}^{\prime}_{\varphi},\widetilde{\mathcal{D}}^{\prime}_{\varphi},\widetilde{\mathcal{L}}^{\prime}_{\varphi})\rightarrow\mathbb{A}^{1}$, whose general fiber is still deformation equivalent to $(\widetilde{Y},\widetilde{D},\widetilde{L})$, but with now the line bundle ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ being relatively big and nef. By [34, Theorem 4.18], a big enough power of ~$\widetilde{\mathcal{L}}^{\prime}_{\varphi}$ is base point free, and so defines a relatively birational contraction
@@
$(\mathcal{Y}_{\varphi},\mathcal{D}_{\varphi},\mathcal{L}_{\varphi})\longrightarrow\mathbb{A}^{1}$ (2.8)